# A Surrogate-Informed Framework for Sparse Grid Interpolation

### v1

@@ 1 Introduction
For instance, stochastic-collocation methods provide a polynomial approximation of a quantity of interest as a function of parameters characterized by probability density functions (PDFs). However, when the number of uncertain parameters is large, the computational cost of standard tensor-product collocation methods grows exponentially, becoming prohibitive for practical applications. In this context, sparse grids have emerged as a powerful and widely adopted technique to address this so-called curse of dimensionality (bungartz2004sparse). Originally introduced by smolyak1963quadrature, sparse grids are built through a hierarchical combination of lower-dimensional tensor product grids, selectively discarding grid points that contribute least to the accuracy of the approximation. This construction dramatically reduces the number of required evaluations while preserving high approximation quality for functions with sufficient regularity. As a result, sparse grids significantly mitigate the curse of dimensionality and make high-dimensional problems more tractable.
However, standard sparse grids are inherently isotropic; they refine the approximation space uniformly across all dimensions and regions of the domain. This approach is suboptimal for a large class of problems characterized by anisotropic or localized features, where the function exhibits different behavior or importance across different input parameters or in specific subregions. This limitation has motivated the development of adaptive sparse grid methods, which aim to allocate computational resources more efficiently by refining the grid non-uniformly (gerstner2003dimension). A common strategy involves using the hierarchical surplus (ma2009adaptive), defined as the difference in value between a new grid point and the interpolation from the previous coarser grid, as an a posteriori error indicator to guide the refinement process. In this context, significant contributions have been made to develop both dimension-adaptive and spatially-adaptive schemes.
@@
The proposed strategy has been implemented in a tool developed by piazzola2022sparse, which constructs a sparse grid interpolant using the Smolyak combination technique, where the final surrogate model is a weighted linear combination of simpler, anisotropic tensor-product grids. The selection of these component grids is governed by an isotropic total degree rule, which systematically determines the multi-indices and corresponding coefficients for the combination formula. To ensure numerical stability and efficiency, each tensor grid is built upon Clenshaw-Curtis nodes arranged according to a nested doubling rule, which guarantees that points evaluated at coarser levels are reused in finer ones. The final approximation at any given point is then obtained by summing the values of the individual Lagrange interpolants from each grid in the combination, weighted by their specific Smolyak coefficient.
@@ 3.1 Validation on analytical benchmark functions
The chosen benchmark functions, Ishigami function [ishigami1990importance], Sobol’s G-function and an obscillating function, test the algorithm’s ability to handle strong non-linearities, parameter interactions, and non-smooth features [saltelli2010variance].
@@ 3.2 Application to the hydrogen-fueled perforated burners case study
The physical model considers a representative unit cell with a single slit, defined by four geometric parameters: slit length ($L$), slit width ($W$), inter-slit spacing ($D$), and plate thickness ($H$), as shown in Figure 17. The simulations describe a hydrogen-air mixture at equivalence ratio $\phi=0.6$ with uniform inlet conditions ($T_{u}=$300\text{\,}\mathrm{K}$\$) and ambient outlet pressure ($p=$1\text{\,}\mathrm{a}\mathrm{t}\mathrm{m}$\$). The solid domain is modeled as stainless steel with thermally-coupled interfaces via Conjugate Heat Transfer. Detailed chemistry is captured through a 9-species, 22-reaction mechanism kee. A steady-state approach is adopted to determine the critical flashback velocity $V_{\mathrm{FB}}$, defined as the cold-flow bulk velocity at the slit entry:
@@
where $\psi=A_{\mathrm{slit}}/A_{\mathrm{tot}}$ is the porosity of the slit. This definition provides an intrinsic measure of the slit’s resistance to flashback, independent of the trivial scaling with $\psi$. Under all investigated conditions, the jet Reynolds number remains within the laminar regime. For a comprehensive description of the numerical model, validation studies, and computational setup, the reader is referred to FRUZZA2023; fruzzaUQ; FRUZZA_PROCI; FRUZZA2025_CF.
Figure 17: Left panel: Slit pattern on the burner plate, highlighting the section of the computational domain. Center panel: Slit geometry (boundary conditions in red) and solid plate mesh. Right panel: Computational domain (fluid zone in light blue, solid zone in gray, boundary conditions in red). Reproduced from FRUZZA2025_CF.
In a previous study Fruzzasub2025, a surrogate-informed sparse-grid methodology was developed to map critical flashback velocities and maximum burner temperatures across the four-dimensional design space. Since each sparse grid point required a costly three-dimensional simulation, an initial sparse grid at level $w=2$ was constructed, combining high-fidelity simulations with surrogate modeling to reduce computational cost. However, the response surfaces generated at this level were found to lack sufficient accuracy, particularly in regions of high sensitivity ($L\approx 0$). To address this limitation, the grid was manually refined by evaluating 33 additional simulated points at level $w=3$ near $L=0$, while the remaining points were populated using interpolated values from the initial surrogate model. Although this physics-informed manual refinement improved accuracy, it relied on prior domain knowledge and lacked a systematic criterion for point selection.
@@
The two scalar output quantities of interest are flashback velocity and maximum burner temperature. Thus, the refinement process can be guided by either quantity. We observed that the resulting performance exhibits no significant differences depending on the variable chosen to drive the refinement, so in what follows, we refer to the process using the flashback velocity as the guiding quantity. Figure 18 presents the discrepancy criterion computed for candidate nodes at level $w+1$, ranked in decreasing order of importance. To enable a direct comparison between the manually refined physics-based approach from Fruzzasub2025 and the surrogate-informed hierarchical strategy, the budget in this study was fixed at 33 points, matching the number of simulations used during the previous empirical refinement phase.

### v2

@@ 1 Introduction
For instance, stochastic-collocation methods provide a polynomial approximation of a quantity of interest as a function of parameters characterized by probability density functions (PDFs). However, when the number of uncertain parameters is large, the computational cost of standard tensor-product collocation methods grows exponentially, becoming prohibitive for practical applications. In this context, sparse grids have emerged as a powerful and widely adopted technique to address this so-called curse of dimensionality (Bungartz and Griebel, 2004). Originally introduced by Smolyak (1963), sparse grids are built through a hierarchical combination of lower-dimensional tensor product grids, selectively discarding grid points that contribute least to the accuracy of the approximation. This construction dramatically reduces the number of required evaluations while preserving high approximation quality for functions with sufficient regularity. As a result, sparse grids significantly mitigate the curse of dimensionality and make high-dimensional problems more tractable.
However, standard sparse grids are inherently isotropic; they refine the approximation space uniformly across all dimensions and regions of the domain. This approach is suboptimal for a large class of problems characterized by anisotropic or localized features, where the function exhibits different behavior or importance across different input parameters or in specific subregions. This limitation has motivated the development of adaptive sparse grid methods, which aim to allocate computational resources more efficiently by refining the grid non-uniformly (Gerstner and Griebel, 2003). A common strategy involves using the hierarchical surplus (Ma and Zabaras, 2009), defined as the difference in value between a new grid point and the interpolation from the previous coarser grid, as an a posteriori error indicator to guide the refinement process. In this context, significant contributions have been made to develop both dimension-adaptive and spatially-adaptive schemes.
@@
The proposed strategy has been implemented in a tool developed by Piazzola and Tamellini (2022), which constructs a sparse grid interpolant using the Smolyak combination technique, where the final surrogate model is a weighted linear combination of simpler, anisotropic tensor-product grids. The selection of these component grids is governed by an isotropic total degree rule, which systematically determines the multi-indices and corresponding coefficients for the combination formula. To ensure numerical stability and efficiency, each tensor grid is built upon Clenshaw-Curtis nodes arranged according to a nested doubling rule, which guarantees that points evaluated at coarser levels are reused in finer ones. The final approximation at any given point is then obtained by summing the values of the individual Lagrange interpolants from each grid in the combination, weighted by their specific Smolyak coefficient.
@@ 3.1 Validation on analytical benchmark functions
The chosen benchmark functions, Ishigami function [Ishigami and Homma (1990)], Sobol’s G-function and an obscillating function, test the algorithm’s ability to handle strong non-linearities, parameter interactions, and non-smooth features [Saltelli et al. (2010)].
@@ 3.2 Application to the hydrogen-fueled perforated burners case study
The physical model considers a representative unit cell with a single slit, defined by four geometric parameters: slit length ($L$), slit width ($W$), inter-slit spacing ($D$), and plate thickness ($H$), as shown in Figure 17. The simulations describe a hydrogen-air mixture at equivalence ratio $\phi=0.6$ with uniform inlet conditions ($T_{u}=$300\text{\,}\mathrm{K}$\$) and ambient outlet pressure ($p=$1\text{\,}\mathrm{a}\mathrm{t}\mathrm{m}$\$). The solid domain is modeled as stainless steel with thermally-coupled interfaces via Conjugate Heat Transfer. Detailed chemistry is captured through a 9-species, 22-reaction mechanism Bilger et al. (1990). A steady-state approach is adopted to determine the critical flashback velocity $V_{\mathrm{FB}}$, defined as the cold-flow bulk velocity at the slit entry:
@@
where $\psi=A_{\mathrm{slit}}/A_{\mathrm{tot}}$ is the porosity of the slit. This definition provides an intrinsic measure of the slit’s resistance to flashback, independent of the trivial scaling with $\psi$. Under all investigated conditions, the jet Reynolds number remains within the laminar regime. For a comprehensive description of the numerical model, validation studies, and computational setup, the reader is referred to Fruzza et al. (2023, 2024b, 2024a, 2025a).
Figure 17: Left panel: Slit pattern on the burner plate, highlighting the section of the computational domain. Center panel: Slit geometry (boundary conditions in red) and solid plate mesh. Right panel: Computational domain (fluid zone in light blue, solid zone in gray, boundary conditions in red). Reproduced from Fruzza et al. (2025a).
In a previous study Fruzza et al. (2025b), a surrogate-informed sparse-grid methodology was developed to map critical flashback velocities and maximum burner temperatures across the four-dimensional design space. Since each sparse grid point required a costly three-dimensional simulation, an initial sparse grid at level $w=2$ was constructed, combining high-fidelity simulations with surrogate modeling to reduce computational cost. However, the response surfaces generated at this level were found to lack sufficient accuracy, particularly in regions of high sensitivity ($L\approx 0$). To address this limitation, the grid was manually refined by evaluating 33 additional simulated points at level $w=3$ near $L=0$, while the remaining points were populated using interpolated values from the initial surrogate model. Although this physics-informed manual refinement improved accuracy, it relied on prior domain knowledge and lacked a systematic criterion for point selection.
@@
The two scalar output quantities of interest are flashback velocity and maximum burner temperature. Thus, the refinement process can be guided by either quantity. We observed that the resulting performance exhibits no significant differences depending on the variable chosen to drive the refinement, so in what follows, we refer to the process using the flashback velocity as the guiding quantity. Figure 18 presents the discrepancy criterion computed for candidate nodes at level $w+1$, ranked in decreasing order of importance. To enable a direct comparison between the manually refined physics-based approach from Fruzza et al. (2025b) and the surrogate-informed hierarchical strategy, the budget in this study was fixed at 33 points, matching the number of simulations used during the previous empirical refinement phase.
@@ 4 Conclusions
5 Acknowledgments
L. Tamellini has been partially supported by the project 202222PACR “Numerical approximation of uncertainty quantification problems for PDEs by multi-fidelity methods (UQ-FLY)” and by the ICSC—Centro Nazionale di Ricerca in High Performance Computing, Big Data, and Quantum Computing, both funded by European Union—NextGenerationEU. Lorenzo Tamellini is member of the Gruppo Nazionale Calcolo Scientifico-Istituto Nazionale di Alta Matematica (GNCS-INdAM)